ISOTOPOLOGUE-SENSITIVE DETECTION USING CHIRPED-PULSE FT-MW SPECTROSCOPY: MI-NOR SPECIES OF PROPOFOL

\underline{ALBERTO LESARRI}, Departamento de Química Física y Química Inorgánica, Facultad de Ciencias, Universidad de Valladolid, E-47011 Valladolid, Spain; JUSTIN NEILL, MATT MUCKLE, STEVEN T. SHIPMAN, BROOKS H. PATE and RICHARD D. SUENRAM, Department of Chemistry, University of Virginia, Charlottesville, VA 22904, USA; WALTHER CAMINATI, Dipartimento di Chimica ”G. Ciamician”, Università di Bologna, I-40126 Bologna, Italy.

The capabilities of chirped-pulse FT-microwave spectroscopy to achieve full-bandwidth (11 GHz) isotopologue-sensitive detection have been tested on the 13-heavy atoms molecule of propofol (2,6-diisopropylphenol). The analysis of the rotational spectrum using moderate signal averaging (10 k FIDs) had previously detected the presence of two conformers arising from the combined internal rotations of the hydroxyl and the two isopropyl groups^{a}. In the new experiment reported here 600 k FIDs were coherently averaged, using three pulsed nozzle sources and reading multiple FIDs per sample injection cycle to reduce the total acquisition time and sample consumption. The new spectrum revealed a very large number of weak transitions, suggesting that full-band ^{13}C sensitivity had been surpassed. The new data have resulted in the assignment of a third conformer of propofol, followed by all twelve ^{13}C-monosubstituted species in natural abundance for the most stable conformer. The isotopic information confirmed the molecular structure for the preferred conformation of propofol, validating the ab initio predictions for this compound. The potential function for the OH internal rotation has been determined using a flexible model.

aA. Lesarri, S. T. Shipman, G. G. Brown, L. Alvarez-Valtierra, R. D. Suenram and B. H. Pate, 63^{rd} OSU Int. Symp. On Mol. Spectrosc., Columbus, OH, 2008, RH07